OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. x intercept = (6, 0); y intercept = (0, -8) x intercept = (-6, 0); y intercept = (0, 8) x intercept = (-8, 0); y intercept = (0, -6) x intercept = (8, 0); y intercept = (0, 6)

OpenStudy (anonymous):

When in Standard Form: Ax + By = C slope = -A/B y-int = C/B

OpenStudy (anonymous):

x-int = C/A

OpenStudy (anonymous):

x-int = 24/4 = ... (ans,0) y-int = 24/-3 = ...(0,ans)

OpenStudy (anonymous):

Don't forget to mark a Best Response. Thanks and take care.